# PrivTuner with Homomorphic Encryption and LoRA: A P3EFT Scheme for Privacy-Preserving Parameter-Efficient Fine-Tuning of AI Foundation Models

### v1

@@ 7            Update auxiliary variables $\bm{z}^{(i)}$, where:
The optimal solution $(p_{n}^{*},B_{n}^{*})$ of problem $\mathbb{P}_{3}$ can be computed by suitably defined functions $\overline{B}_{n}(\cdot,\cdot)$ and $\overline{p}_{n}(\cdot,\cdot)$ such that:
@@ VIII Experiments
In this section on experiments, Sec. VIII-A and Sec. VIII-B provide the curve fitting results and parameter settings, respectively. Sec. VIII-C compares our algorithm’s optimization performance with benchmarks, and Sec. VIII-D investigates the influence of the weight parameter.
VIII-A Curve Fitting Results
We employ the CKKS scheme to test the number of CPU cycles required for different operations: encryption, addition and multiplication. It is also worth noting that CKKS is not the only option, and our analysis can be adapted to other FHE mechanisms. Specifically, we measure the number of required CPU cycles by obtaining the average running time and computation frequency on a single NVIDIA GeForce RTX 2080 Ti. The curve fitting results are shown in Fig. 3(a). Next, we discuss the curve fitting of the privacy protection level. To more comprehensively measure the security level under different $\lambda$, we run the LWE-estimator under three attacks and give the results in Table I. The minimum one of three attacks is adopted as the corresponding privacy protection. Additionally, Fig. 3(b) presents the curve fitting performance.
(a) CPU cycles for each Encryption (Enc), Addition (Add) and Multiplication (Mul) under different $\lambda$ and the fitting performance.
(b) Privacy protection level measured by the LWE-estimator under different $\lambda$ from Table I and the fitting performance.
@@ VIII-A Curve Fitting Results
TABLE I: Estimated running times (in bits) for three attacks under different $\lambda$ with $q=210$ in CKKS.
@@
2048 NaN 44.5 46.3
4096 66.0 65.8 66.9
6144 97.2 96.9 98.1
8192 131.0 130.8 131.9
12288 205.1 204.8 205.9
16384 285.4 281.7 286.2
24576 459 458.7 459.5
32768 645.0 NaN NaN
The fitted functions and corresponding constants obtained from the curve fitting are listed in Table II.
TABLE II: Result of Fitted Functions.
@@
Encryption $C_{1}(\lambda+C_{2})^{2}$, where $C_{1}=0.0066$, $C_{2}=22421$
Addition $C_{3}\lambda+C_{4}$, where $C_{3}=29.0862$, $C_{4}=-30253$
Multiplication $C_{5}\lambda+C_{6}$, where $C_{5}=68.6450$, $C_{6}=57511$
Privacy protection level $C_{7}\lambda+C_{8}$, where $C_{7}=0.0205$, $C_{8}=-30.73$
VIII-B Parameter Settings
We provide a detailed description of the experiments’ default parameter settings. The total number of mobile devices $N$ is $10$. The number of polynomial degree options $M$ is $3$ and we have $\{\lambda_{\textnormal{o}1},\lambda_{\textnormal{o}2},\lambda_{\textnormal{o}3}\}=\{2^{12},2^{13},2^{14}\}$. The number of samples $D_{n}$ is 10, and the number of encrypted parameters per sample $s_{n}$ is $10^{4}$. The number of additions $a_{n}$ and multiplications $m_{n}$ required for prediction are set to $10^{5}$ and $10^{6}$. The number of CPU cycles $C_{n}^{\textnormal{other}}$ for other operations is set as $10^{9}$. The size of encrypted data for transmission $d_{n}$ is $3\times 10^{9}$ bits. The privacy concern $c_{n}$ of each device is randomly selected from $\{10,5,1\}$, and the weight parameter $\omega$ for the overall privacy level is set to $1$ (unless configured otherwise).
Then, we introduce settings regarding computation and communication resources. The total server computation capacity $f^{\textnormal{total}}$ is $10$ GHz. The maximum device frequency $g_{n}^{\textnormal{max}}$ is $3$ GHz. The path loss model is $128.1+37.6\log(\text{distance})$, with an $8$ dB standard deviation for shadow fading, and the unit of distance is kilometer. The noise power spectral density $N_{0}$ is $-174$ dBm/Hz. The effective switched capacitance $\kappa$ is $10^{-28}$. The total available bandwidth $B^{\textnormal{total}}$ is $10$ MHz. The maximum device transmission power $p_{n}^{\textnormal{max}}$ is $20$ dBm. The maximum time budget on devices $T^{\textnormal{max}}_{D}$ is $4000$ s, and the time budget on the server $T^{\textnormal{max}}_{S}$ is set as $5000$ s (unless configured otherwise).
VIII-C Comparison with Benchmarks on Energy Optimization
@@
    Average Allocation: In this case, we set up the resources of the server and mobile devices fairly. Specifically, we set each $f_{n},g_{n},p_{n},B_{n}$ as $f^{\textnormal{total}}/N,g_{n}^{\textnormal{max}}/2,p_{n}^{\textnormal{max}}/2,B^{\textnormal{total}}/N$, respectively. Besdies, the choices of $\bm{\lambda}$ are all set to $\lambda_{\textnormal{o}1}$.
@@
Total server frequency. Here we configure the total server CPU frequency $f^{\textnormal{total}}$ from $5$ GHz to $9$ GHz. Since the optimal frequencies need to be minimized to bring lower energy, we can see from Fig. 4(c) that curves of “Optimize $\bm{f}$, $\bm{g}$ only” and our proposed method do not change as $f^{\textnormal{total}}$ increases.
@@
Time budget on devices. Here we configure the time budget on devices $T_{D}^{\textnormal{max}}$ from $3000$ s to $5000$ s. A higher $T_{D}^{\textnormal{max}}$ allows mobile devices to adopt lower computation frequencies and transmission powers to further decrease energy consumption. Hence, a slower downward trend can be observed in Fig. 4(e).
Time budget on server. Here we vary the time budget on server $T_{S}^{\textnormal{max}}$ from $3000$ s to $7000$ s. A higher $T_{S}^{\textnormal{max}}$ allows the server to allocate lower computation frequencies $\bm{f}$ to mobile devices, thus decreasing energy consumption. Thereafter, the energy optimization of our proposed method and “Optimize $\bm{f}$, $\bm{g}$ only” decreases as $T_{S}^{\textnormal{max}}$ grows.
VIII-D Impact of weight parameter
Now we check how the weight parameter $\omega$ affects the tradeoff between energy and overall privacy level. We also detail the choices of $\bm{\lambda}$ by devices with different levels of privacy protection $c_{n}$. Specifically, we raise time budgets $T^{\textnormal{max}}_{D}$ and $T^{\textnormal{max}}_{S}$ to $10000$ s, and we adjust $\omega$ from $1$ to $10$. The experimental results are reported in Fig. 5(a) and (b).
@@ VIII-D Impact of weight parameter
IX Conclusion
@@
In summary, this paper introduced a novel scheme named PrivTuner for P3EFT tasks of AI foundation models. Specifically, PrivTuner integrates LoRA with FHE schemes, allowing the model owner server and external mobile devices to collaboratively implement secure and efficient fine-tuning. Additionally, we also introduce a resource allocation method to jointly optimize energy consumption and overall privacy level within PrivTuner. A theoretical analysis of time complexity, solution quality and convergence is provided. The experimental results demonstrate the effectiveness and superiority of our approach, making it a promising solution to privacy and resource problems in fine-tuning foundation models.
Appendix A: More details of Algorithm 2
@@
$\mathcal{L}_{1}(\bm{\lambda},\alpha)\!=\!$ $\sum_{n=1}^{N}\!\kappa y_{1}(\lambda_{n})M_{n}\overline{f}_{n}^{2}(\lambda_{n})\!+\!\sum_{n=1}^{N}R_{l}\kappa y_{2}(\lambda_{n})D_{n}\overline{g}_{n}^{2}(\lambda_{n})$
$-\omega\sum_{n=1}^{N}c_{n}y_{5}(\lambda_{n})+\alpha(\sum_{n=1}^{N}\overline{f}_{n}(\lambda_{n})-f^{\textnormal{total}}).$ (39)
Here we first give the KKT conditions of (Appendix A: More details of Algorithm 2):
$\frac{\partial\mathcal{L}_{1}}{\partial\lambda_{n}}=$ $\frac{6\kappa M_{n}^{3}C_{1}^{3}(\lambda_{n}+C_{2})^{5}}{(T^{\textnormal{max}}_{\textnormal{en}})^{2}}+\frac{3\kappa R_{l}^{3}D_{n}^{3}(C_{3}a_{n}\!+\!C_{5}m_{n})^{3}\lambda_{n}^{2}}{(T^{\textnormal{max}}_{\textnormal{ft}})^{2}}$
$\!-\!\omega c_{n}C_{7}\lambda_{n}\!+\!\alpha\frac{2C_{1}(\lambda_{n}\!+\!C_{2})M_{n}}{T^{\textnormal{max}}_{\textnormal{en}}}\!=\!0,~{}\forall n\in\mathcal{N}.$ (40)
@@
    $\sum_{n=1}^{N}\overline{f}_{n}(\max\{\min\{\lambda_{\textnormal{o}m},\!\lambda_{n}^{\textnormal{max}},\!\overline{\lambda}_{n}(\alpha)\},\!\lambda_{o1}\})\!=\!f^{\textnormal{total}}.$ (45)
@@
[…]
@@
Appendix B: Proof of Theorem 1

### v3

@@ 7    Update auxiliary variables $\bm{z}^{(i)}$, where:
The optimal solution $(p_{n}^{*},B_{n}^{*})$ of problem $\mathbb{P}_{4}$ can be computed by suitably defined functions $\overline{B}_{n}(\cdot,\cdot)$ and $\overline{p}_{n}(\cdot,\cdot)$ such that:
@@ VIII Experiments
VIII-A Macrobenchmark
We implement the CKKS-supported BERT-Tiny in C++ and Python, utilizing the OpenFHE [72] library. Specifically, we evaluate the experiments on a Linux x86_64 machine equipped with Intel(R) Xeon(R) Gold 5218R CPU@2.10GHz. In addition, we also use HEXL [73] to accelerate OpenFHE. Following Rovida et al. [29], we configure the polynomial degree $\lambda=2^{15}$, coefficient modulus $q=1767$ bits, with a precision factor of $55$ bits. We evaluated the model accuracy on three GLUE datasets: SST-2, MRPC and RTE. The batch size is set as $32$; the learning rate is $1\times 10^{-5}$, and the training epoch is $10$. The accuracy performance is reported in Table I. Besides, we also provide a computation runtime breakdown when doing inference on one sample of $8$ tokens from the SST-2 dataset, which is as shown in Table II.
TABLE I: Test performance on different GLUE datasets.
Dataset BERT-Tiny FHE-BERT-Tiny Loss $\downarrow$
SST-2 (Acc.) 0.823 0.790 0.033
MRPC (Acc.) 0.703 0.675 0.028
WNLI (Acc.) 0.601 0.564 0.037
CoLA (Acc., M Corr.) 0.691, 0 0.691, 0 0.000, 0
TABLE II: A computation runtime breakdown for inference.
Operation Time Consumption (s)
Encryption (client) 0.7106
Prediction (server) 163.3211
Decryption (client) 0.0119
Total 164.0436
Besides, since we aim to optimize the polynomial degree $\lambda$ in CKKS to balance the overhead and security, we also report the runtime and security level under different $\lambda$. From Table III, we can see that a higher $\lambda$ increases the runtime obviously, but also benefits the security level. We also want to emphasize that even if the overall runtime is often hundreds of seconds, the primary bottleneck lies in the server-side encrypted predictions. In contrast, client-side tasks are limited to encryption and decryption, keeping the computational demands relatively manageable for mobile devices (Table II).
TABLE III: Runtime and security level under different $\lambda$.
Method $\lambda$ Runtime Security Level
FHE-BERT-Tiny $2^{15}$ 164.04 s 66.1
$2^{16}$ 330.13 s 128.4
$2^{17}$ 719.64 s 277.0
PEFT (LoRA) - 0.067 s 0 (unprotected)
VIII-B Curve Fitting
Here, we test the number of CPU cycles required for different operations: encryption, addition, multiplication and rotation. It is also worth noting that CKKS is not the only option, and our analysis can be adapted to other FHE mechanisms. Specifically, we measure the number of required CPU cycles by obtaining the average running time and computation frequency. The curve fitting results are shown in Fig. 3(a). To fit the privacy protection level, we run the LWE-estimator under three attacks and give the results in Table IV. The minimum one of three attacks is adopted as the corresponding privacy protection. Fig. 3(b) presents the curve fitting performance.
(a) The CPU cycles for each Addition (Add), Encryption (Enc), Multiplication (Mul) and Rotation (Rot) under different $\lambda$.
(b) Privacy protection level measured by the LWE-estimator under different $\lambda$ from Table IV and the fitting performance.
@@ VIII-B Curve Fitting
TABLE IV: Estimated running times (in bits) for three attacks under different $\lambda$ with $q=1767$ bits in CKKS.
@@
16384 47.6 47.6 48.6
24576 50.5 50.3 50.9
32768 66.3 66.1 66.6
49152 96.1 96.1 96.7
65536 128.5 128.4 129.4
98304 199.7 199.5 200.7
131072 277.0 296.6 278.1
The fitted functions and corresponding constants obtained from the curve fitting are listed in Table V.
TABLE V: Result of fitted functions.
@@
Encryption $C_{0}(\lambda+C_{1})^{2}$, where $C_{0}=0.012$, $C_{1}=6.45\times{10}^{4}$
Addition $C_{2}\lambda+C_{3}$, where $C_{2}=165.15$, $C_{3}=-6.34\times{10}^{5}$
Multiplication $C_{4}\lambda+C_{5}$, where $C_{4}=5282.55$, $C_{5}=-1.91\times{10}^{7}$
Rotation $C_{6}\lambda+C_{7}$, where $C_{6}=4979.69$, $C_{7}=-3.19\times{10}^{7}$
Security level $C_{8}\lambda+C_{9}$, where $C_{8}=0.0020$, $C_{9}=1.4789$
VIII-C Parameter Settings
We provide a detailed description of the experiments’ default parameter settings. The total number of mobile devices $N$ is $10$. The number of polynomial degree options $M$ is $3$ and we have $\{\lambda_{\textnormal{o}1},\lambda_{\textnormal{o}2},\lambda_{\textnormal{o}3}\}=\{2^{15},2^{16},2^{17}\}$. The number of samples $D_{n}$ is 16, and the number of tokens $s_{n}$ for encryption is 10. The estimated number of additions $a_{n}$, multiplications $m_{n}$, and rotations $o_{n}$ required for prediction are set to $1660$, $260$ and $1460$ (Please see Appendix B for details).
The number of CPU cycles $C_{n}^{\textnormal{other}}$ for other operations is set as $10^{9}$. The size of encrypted data for transmission $d_{n}$ is $3\times 10^{9}$ bits. The privacy concern $c_{n}$ of each device is randomly selected from $\{10,5,1\}$, and the weight parameter $\omega$ for the overall privacy level is set to $1$ (unless configured otherwise).
Then, we introduce settings regarding computation and communication resources. The total server computation capacity $f^{\textnormal{total}}$ is $20$ GHz. The maximum device frequency $g_{n}^{\textnormal{max}}$ is $3$ GHz. The path loss model is $128.1+37.6\log(\text{distance})$, with an $8$ dB standard deviation for shadow fading, and the unit of distance is kilometer. The noise power spectral density $N_{0}$ is 174-174 dBm/Hz. The effective switched capacitance $\kappa$ is $10^{-28}$. The total available band width $B^{\textnormal{total}}$ is $10$ MHz. The maximum device transmission power $p_{n}^{\textnormal{max}}$ is $20$ dBm. The maximum time budget on devices $T^{\textnormal{max}}_{D}$ is $1000$ s, and the time budget on the server $T^{\textnormal{max}}_{S}$ is set as $3000$ s (unless configured otherwise).
VIII-D Comparison with Benchmarks on Energy Optimization
@@
    Average Allocation: In this case, we set up the resources of the server and mobile devices fairly. Specifically, we set each $f_{n},g_{n},p_{n},B_{n}$ as $f^{\textnormal{total}}/N,g_{n}^{\textnormal{max}}/2,p_{n}^{\textnormal{max}}/2,B^{\textnormal{total}}/N$, respectively. Besides, the choices of $\bm{\lambda}$ are all set to $\lambda_{\textnormal{o}1}$.
@@
Total Server Frequency. Here we vary the total server CPU frequency $f^{\textnormal{total}}$ from $10$ GHz to $14$ GHz. Since the optimal frequencies need to be minimized to bring lower energy, we can see from Fig. 4(c) that curves of “Optimize $\bm{f}$, $\bm{g}$ only” and our proposed method do not change as $f^{\textnormal{total}}$ increases.
@@
Time Budget on Devices. Here we configure the time budget on devices $T_{D}^{\textnormal{max}}$ from $1000$ s to $3000$ s. A higher $T_{D}^{\textnormal{max}}$ allows mobile devices to adopt lower computation frequencies and transmission powers to further decrease energy consumption. Hence, a slower downward trend can be observed in Fig. 4(e).
Time Budget on Server. Here we vary the time budget on server $T_{S}^{\textnormal{max}}$ from $2000$ s to $6000$ s. A higher $T_{S}^{\textnormal{max}}$ allows the server to allocate lower computation frequencies $\bm{f}$ to mobile devices, thus decreasing energy consumption. Thereafter, the energy optimization of our proposed method and “Optimize $\bm{f}$, $\bm{g}$ only” decreases as $T_{S}^{\textnormal{max}}$ grows.
VIII-E Impact of Weight Parameter on Energy and Privacy
Now we check how the weight parameter $\omega$ affects the tradeoff between energy and overall privacy level. We also detail the choices of $\bm{\lambda}$ by devices with different levels of privacy protection $c_{n}$. Specifically, we raise time budgets $T^{\textnormal{max}}_{D}$ and $T^{\textnormal{max}}_{S}$ to $1500$ s and $5000$ s, and we adjust $\omega$ from $1$ to $10$. The experimental results are reported in Fig. 5(a) and (b).
@@ VIII-E Impact of Weight Parameter on Energy and Privacy
IX Discussion & Conclusion
Privacy of Adapters. One potential concern for PrivTuner is that the adapter may leak some information about data. We use unencrypted adapters to facilitate the following inference services after fine-tuning. Given this limitation, we adopt an honest but curious security model, where the server is assumed to refrain from malicious actions (e.g., data reconstruction attacks). Besides, adapters are much smaller than the foundation model and contain far less information. We also consider potential solutions in future work, including fine-tuning in a trustworthy environment [74, 75], or using MPC techniques [76, 77] to distribute updates across multiple servers without any single server having full adapter information.
Malicious Security. Supporting malicious security poses challenges for our proposed PrivTuner. If either the server or the client maliciously deviates from the protocol (e.g., by intentionally sending incorrect data), then the fine-tuning process may be compromised. However, even under a malicious adversary model, PrivTuner maintains security by isolating client and server roles. A malicious client gains no extra model information and only risks its own process if deviating, while a malicious server cannot deduce true labels from plaintext loss due to encrypted predictions. Additionally, client-specific FHE secret keys prevent the server from gaining further information through collusion. In future work, we will explore how to adapt advanced strategies from current malicious secure studies [78, 79] to further enhance the robustness of PrivTuner.
Accelerate FHE with Hardware and GPU Backends. While our implementation leverages HEXL to accelerate FHE computations, the overhead remains challenging for real-time applications. GPU-based FHE acceleration [28] has shown promise, with emerging libraries [80, 81] offering support. However, these are still evolving and seldom used for complex neural networks. As our framework is higher-level and network-oriented, it can seamlessly integrate future GPU-accelerated solutions, enhancing both efficiency and security.
@@ IX Discussion & Conclusion
Conclusion. In summary, this paper introduced a novel scheme named PrivTuner for P3EFT tasks of AI foundation models. Specifically, PrivTuner integrates LoRA with FHE schemes, allowing the model owner server and external mobile devices to collaboratively implement secure and efficient fine-tuning. Additionally, we also introduce a resource allocation method to jointly optimize energy consumption and overall privacy level within PrivTuner. A theoretical analysis of time complexity, solution quality and convergence is provided. The experimental results demonstrate the effectiveness and superiority of our approach, making it a promising solution to privacy and resource problems in fine-tuning foundation models.
Appendix A Approximating Non-linear Functions
Here, we briefly present the method of Rovida et al. [29] to approximate the non-linear functions in BERT-Tiny.
Softmax. Given an input vector $\bm{x}\in\mathbb{R}^{d}$, the Softmax function can be computed as:
$\mathrm{Softmax}(\bm{x})=\left[\frac{\exp({x_{i}})}{\sum_{j=1}^{d}\exp({x_{j}})}\biggm|_{i\in[d]}\right].$ (39)
The main challenge is to efficiently calculate the underlying exponential function $\mathrm{exp}(x)$ and division computation $1/x$. The exponentiation is approximated using the Maclaurin series:
$\mathrm{exp}(x)\approx\sum_{i=0}^{n}\frac{x^{i}}{i!}.$ (40)
By doing so, we can limit the average error within $1\times 10^{-7}$ with the degree $n=6$, as shown in Table VI.
For the division operation $1/x$, we utilize a Chebyshev polynomial supported by OpenFHE to approximate it as follows:
$1/x=\frac{c_{0}}{2}+\sum_{i=1}^{n}c_{i}T_{i}(x),$ (41)
where $n$ is the polynomial degree, $T_{i}(x)=\cos{(i\arccos{x})}$ provide an orthogonal basis of polynomials on the interval [1,1][-1,1] with the weight function $1/\sqrt{1-x^{2}}$, and $c_{i}$ is the coefficient for fitting. Following [29], we set $n=119$ and the input interval for fitting as $[2,5000]$. By doing so, we can observe that the average error can be limited within $1\times 10^{-4}$ with a multiplicative depth of $8$ in Table VI.
GeLU. The GeLU (Gaussian Error Linear Unit) activation function is defined as follows:
$\mathrm{GeLU}(x)=\frac{x}{2}\big(1+\mathrm{erf}(\frac{x}{\sqrt{2}})\big),$ (42)
where $\mathrm{erf}(\cdot)$ denotes the Gaussian error function which is given by $\mathrm{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$. In particular, we approximate the whole GELU function utilizing a Chebyshev polynomial:
$\mathrm{GeLU}(x)=\frac{c_{0}}{2}+\sum_{i=1}^{n}c_{i}T_{i}(x),$ (43)
where we set $n=59$, and the average error on interval of [18,8][-18,8] could be limited within $4\times 10^{-8}$, as shown in Table VI.
LayerNorm. For a given vector $\bm{x}\in\mathbb{R}^{d}$, the LayerNorm function is defined as follows:
$\mathrm{LayerNorm}(\bm{x})=\Big[\frac{(x_{i}-\mu)}{\sigma}\cdot\gamma+\beta\mid_{i=1,\ldots,d}\Big],$ (44)
where $\mu=\sum_{i=1}^{d}x_{i}/d$ and $\sigma=\sqrt{\sum_{i=1}^{d}(x_{i}-\mu)^{2}}$ are mean and standard deviation, and $\gamma$ and $\beta$ are affine transform parameters. A precomputed LayerNorm is implemented, where the values of $\mu$ and $\sigma$ are experimentally observed and precomputed. Hence, the LayerNorm function could be simplified as:
$\mathrm{LayerNorm}(\bm{x})\approx\Big[(x_{i}-\hat{\mu})\cdot\hat{\sigma}\cdot\gamma+\beta\mid_{i=1,\ldots,d}\Big],$ (45)
where $\hat{\mu}$ and $\hat{\sigma}$ are the precomputed mean vectors for $\mu$ and $1/\sigma$. By doing so, we could simplify a lot of circuits at the expense of a 0.022 accuracy drop on SST-2, as reported by [29]. For better accuracy performance, we can only precompute the value of $\sigma$, as the computation of $\mu=\sum_{i=1}^{d}x_{i}/d$ is actually linear and easily supported by FHE.
Pooler. After stacked encoders, we implement a pooling layer as the task layer. A fully connected layer is evaluated on the vector of [CLS] token, followed by an activation function of the hyperbolic tangent: $\mathrm{tanh}(x)=\mathrm{sinh}(x)/\mathrm{cosh}(x)$. Similarly, we view $\mathrm{tanh}(x)$ as a whole and use a Chebyshev polynomial to do the approximation. The observed interval of approximation required is [20,20][-20,20], and a $119$-degree polynomial is enough to limit the average error within $3\times 10^{-5}$.
TABLE VI: Non-linear Functions Approximations.
Function Method Interval Degree Avg Error
exp(x) Maclaurin series [-1,1] 6 $3.94\times 10^{-8}$
1/x Chebyshev polynomial [2,5000] 200 $9.18\times 10^{-5}$
GeLU(x) Chebyshev polynomial [-18,8] 59 $3.99\times 10^{-8}$
tanh(x) Chebyshev polynomial [-20,20] 119 $2.51\times 10^{-5}$
Appendix B Count of arithmetic operations for simulation
We provide a rough estimation of the count of FHE operations based on the employed FHE-BERT-Tiny model [29] and its utilized FHE techniques.
Specifically, we first report the count of operations to the basic supporting algorithms [29] (e.g., RotSum) in Table VII. This could be done by checking their released codes at https://github.com/narger-ef/FHE-BERT-Tiny/tree/main. In our experiments, we found the major bulk of linear computation overhead is caused by calling these algorithms for large matrix-matrix multiplications. Thus, we are able to roughly estimate the number of different operations by counting the number of calls to these algorithms. For instance, we could set the number of tokens as $10$, and the numbers of multiplications, additions and rotations in a BERT-Tiny encoder can be calculated as shown in Table VIII. Hence, the number of multiplications, additions and rotations could be set into $260$, $1660$ and $1460$ for a BERT-Tiny model consisting of $2$ encoders. Since the number of operations is considered constant in the optimization problem, it does not affect the solution process.
TABLE VII: The count of operations to basic algorithms, and $L$ denotes the number of tokens for an input.
Algorithm #Multiplications #Additions #Rotations
RotSum 0 77 77
MatMulRE $L$ $8L$ $7L$
MatMulCR $L$ $8L$ $7L$
MatMulRELarge $4L$ $32L$ $31L$
MatMulCRLarge $4L$ $11L$ $7L$
TABLE VIII: The count of operations in an encoder for a sentence with $10$ tokens.
Layer #Multiplications #Additions #Rotations
Self-Attention $50$ $400$ $350$
FFNN $80$ $430$ $380$
Total $130$ $830$ $730$
Appendix C More details of Algorithm 2
@@
$\mathcal{L}_{1}(\bm{\lambda},\alpha)\!=\!$ $\sum_{n=1}^{N}\!\kappa y_{1}(\lambda_{n})s_{n}\overline{f}_{n}^{2}(\lambda_{n})\!+\!\sum_{n=1}^{N}\kappa y_{2}(\lambda_{n})D_{n}\overline{g}_{n}^{2}(\lambda_{n})$
$-\omega\sum_{n=1}^{N}\sigma_{n}y_{6}(\lambda_{n})+\alpha(\sum_{n=1}^{N}\overline{f}_{n}(\lambda_{n})-f^{\textnormal{total}}).$ (47)
Here we first give the KKT conditions of (C):
$\frac{\partial\mathcal{L}_{1}}{\partial\lambda_{n}}=$ $\frac{6\kappa M_{n}^{3}C_{0}^{3}(\lambda_{n}+C_{1})^{5}}{(T^{\textnormal{max}}_{\textnormal{en}})^{2}}\!-\!\omega c_{n}C_{7}\lambda_{n}\!+\!\alpha\frac{2C_{0}(\lambda_{n}\!+\!C_{1})M_{n}}{T^{\textnormal{max}}_{\textnormal{en}}}$
$+\frac{3\kappa D_{n}^{3}(C_{2}a_{n}\!+\!C_{4}m_{n}\!+C_{6}o_{n}\!)^{3}\lambda_{n}^{2}}{(T^{\textnormal{max}}_{\textnormal{ft}})^{2}}\!=\!0,~\forall n\in\mathcal{N}.$ (48)
@@
    $\hskip-5.0pt\sum_{n=1}^{N}\overline{f}_{n}(\max\{\min\{\lambda_{\textnormal{o}m},\!\lambda_{n}^{\textnormal{max}},\!\overline{\lambda}_{n}(\alpha)\},\!\lambda_{o1}\})\!=\!f^{\textnormal{total}}.$ (53)
@@
  • Select a $\widehat{\lambda}_{k}$ such that it is not an integer in $\{\lambda_{\textnormal{o}1},\ldots,\lambda_{\textnormal{o}m}\}$ for separation: $\lambda_{k}^{-}=\lfloor\widehat{\lambda}_{k}\rfloor$ and $\lambda_{k}^{+}=\lceil\widehat{\lambda}_{k}\rceil,$ where $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ denote floor and ceil functions to integers in $\{\lambda_{\textnormal{o}1},\ldots,\lambda_{\textnormal{o}m}\}$, and $\lambda_{k}^{-},\lambda_{k}^{+}$ are rounded results.
@@
Appendix D Proof of Theorem 1